Ms. Spencer started a fashion magazine company in which customers were required to pay her $120 in advance for a year's subscription in which the customer receives one magazine each month. In December 2019 she received 10,000 subscription orders and in January 2020 she published her first magazine. Please prepare the journal entry for the issuance of the magazine in January. Select one: a. Debit: Unearned Subscription Revenue 1,200,000 Credit: Subscription Revenue 1,200,000 b. Debit: Cash 120,000 Credit: Subscription Revenue 120,000 c. Debit: Unearned Subscription Revenue 100,000 Credit: Subscription Revenue 100,000 d. Debit: Unearned Subscription Revenue 120,000 Credit: Subscription Revenue 120,000
Answer #1
Total subscription advance received = 10000*120 = $1200000
Revenue to be recognized per month = 1200000/12 = $100000
The journal entry for the issuance of the magazine in January is:
Debit: Unearned Subscription Revenue 100,000 Credit: Subscription Revenue 100,000
Option C is correct